# Science Out of Its Ivory Tower: Improving Accessibility with Reinforcement Learning

### v1

@@ 4.1 Language Modeling via Proximal Policy Optimization
At the core of our approach is language modeling with Proximal Policy Optimization (PPO) (Schulman et al., 2017) guided by two accessibility measures. A causal language model trained on large corpora can generate the next token based on the current sequence, which is useful in the context of reinforcement learning for developing a policy model that determines the most appropriate next token to maximize the expected return in terms of document readability.
@@
Here, $J(\pi_{\theta})$ represents the expected return when following policy $\pi_{\theta}$. The reward $r(s_{t},a_{t})$ is estimated for each time step $t$ in the trajectory $\tau=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{T-1},a_{T-1})$, where $s_{t}$ is the sequence of tokens at time $t$, formed as the concatenation of $s_{0}$ and the tokens $a_{0},a_{1},\ldots,a_{t-1}$. This formula iteratively computes the rewards given the current sequence $s_{t}$ and the token $a_{t}$ chosen by the policy. The $\beta_{\text{KL}}$-weighted KL divergence term $\text{KL}(\pi_{\theta}(a_{t}\mid s_{t})\parallel\pi_{\theta_{\text{SFT}}}(a_{t}\mid s_{t}))$ is applied at every step of sequence generation to ensure the policy does not deviate significantly from the supervised fine-tuned model. This is crucial because, without such a constraint, the policy model might quickly learn to output whatever the reward model favors to maximize its return, which can lead to undesirable behaviors. For instance, the model might repeatedly output a frequent word (“is is is …”), achieving a high reward based solely on accessibility measures but lacking meaningful content. Following Stiennon et al. (2020), $\beta_{\text{KL}}$ is dynamically adjusted by targeting a specific KL divergence between $\pi_{\theta}$ and $\pi_{\theta_{\text{SFT}}}$ using a capped proportional controller in logarithmic space. The benefit of using a dynamic KL control, as opposed to a fixed one, is that it allows the model to adapt more flexibly to different stages of training, accommodating varying levels of KL divergence between the policy and SFT model. For further details, see Appendix A for the specific PPO implementation we are using and Appendix B for the dynamic proportional controller.
@@ 4.2 Reward Function
The reward function evaluates the overall quality of the output ($s_{T}$). After the initial failures of testing a traditional readability measure (i.e., ARI) as the criterion, we decided to use a balance of two accessibility measures: average sentence length in words and word accessibility, adopted from Riddell and Igarashi (2021), to guide the optimization.
@@
A word’s accessibility is approximated by how frequently it appears in a large reference corpus. We chose the English Wikipedia corpus due to its domain similarity and applied a Moses tokenizer, yielding a vocabulary of 14.6 million types from a total of 3.6 billion tokens. If a token is among the most common 100,000 types, we report its frequency per billion tokens as its accessibility measure. Otherwise, we estimate its frequency using ridge regression with an $\ell 2$-norm coefficient equal to $1.0$. This model allows us to make serviceable estimates of the frequency of arbitrary tokens, including tokens that do not appear in the reference corpus. This model takes as input the token’s length in Unicode code points, its byte unigrams, byte bigrams, and byte trigrams. The model estimates the token’s log frequency per 1 billion tokens. We used the natural logarithm of frequencies per billion tokens as the measure of word accessibility.33endnote: 3We have faithfully followed the experiment of Riddell and Igarashi (2021) with three differences. First, our reference corpus is the English Wikipedia, whereas the original study used the Common Crawl News corpus. Second, we did not discard duplicated sentences as Riddell and Igarashi (2021) did, because we found that sentence duplication is not common in Wikipedia. Third, the original study reported word inaccessibility scores by negating the logarithm of frequency per billion. We report accessibility, without negation, because it is more naturally suited to serve as a reward. Refer to Riddell and Igarashi (2021, pp. 1186–1187) for the training of the ridge regression. For example, the accessibility score for “big” is 11.8, while “colossal” scores 7.3. Despite being comparable in meaning, the model’s production of the latter will receive fewer rewards. Coefficient $\beta_{\text{{wa}}}$ is to control the scale of the credit given for word accessibility.
@@
Sentence length is also determined by a Moses tokenizer, which preserves hyphenation and splits contractions.44endnote: 4For the Moses rule-based tokenizer, we use the sacremoses Python package. We negate the value of sentence length for intuitive calculation of the rewards for optimization.
@@ 5.1 Training
We initialized the policy models $\pi_{\theta}$ by adopting the Gemma-2B checkpoint reported by Wang and Clark (2024) ($\pi_{\theta_{\text{SFT}}}$). The original Gemma-2B was trained on three trillion tokens, consisting of publicly available data as well as proprietary datasets comprising “primarily English data from web documents” (Mesnard et al., 2024). The specific checkpoint we adopted was fine-tuned using the SASS corpus in a straightforward manner. It was chosen for its strong performance in simplification quality, its faithfulness, and its relatively compact size.
@@
The two accessibility rewards were weighted as follows: the word accessibility reward was set to $\beta_{\text{WA}}=4.0$, and we report four models with gradually increasing $\beta_{\text{SL}}$ values of $0.05$, $0.08$, $0.1$, and $0.2$.55endnote: 5Because word accessibility is in logarithmic space, we subtracted 10 from the estimated word accessibility and reset any values lower than 10 to 0 to keep them within a reasonable range. For adaptive control of the per-token semantic reward, we started with an initial $\beta_{\text{KL}}=0.2$ and targeted a KL divergence of $8.0$ nats during the training course, capping it in the range between $0.15$ and $0.25$. We used a micro batch size of $4$, with each sequence used to run the PPO algorithm for $4$ epochs using importance sampling, with gradient accumulation steps set to $4$. We used a clip range of $0.2$ for the policy gradient and value function estimation to ensure stability. The value function coefficient was set to $0.1$. The optimization used standard AdamW optimizer parameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1\times 10^{-8}$) (Kingma and Ba, 2015; Loshchilov and Hutter, 2017). The learning rate was fixed at $1\times 10^{-6}$. The training was conducted on 2 H100 (80GB) GPUs using mixed precision training in bfloat16. The sampling temperature was set to $0.7$ in the rollout phase. Following Huang et al. (2024), we assigned unfinished roll-outs (indicated by the missing end of sequence token) with a fixed low score to encourage the model to generate complete simplified narratives. We also report a model trained as dictated by ARI (i.e., RLARI; described in Section 1.2), with all other parameters kept the same as those guided by the accessibility measures.66endnote: 6We tested other hyperparameters for guiding the optimization of RLARI in pilot studies, but obtained similar results.
@@ 5.2.1 Semantic Retention
BERTScore calculates the cosine similarity between each token in the candidate sentence and each token in the reference sentence using contextual embeddings from a pre-trained language model; its results align well with human judgment on semantic similarity evaluation (Zhang et al., 2019). It is not directly influenced by lexical overlap, making it more suitable for evaluating simplification systems than are metrics that rely on matching words, such as BLEU (Papineni et al., 2002). For our evaluation, we used embeddings from the 18th layer of a BERT-large-uncased model and reported the F1 score. This choice is based on prior findings indicating that the 18th layer yields a strong Pearson correlation (0.72) on the WMT16 To-English benchmark (Zhang et al., 2019).
@@ 5.2.2 Simplification & Accessibility
SARI (System output Against References and against the Input sentence) is specifically designed to evaluate text simplification (Xu et al., 2016). It aims to measure how well a simplified text retains the original meaning while improving readability. SARI provides a balanced measure of how well a text simplification system performs by focusing on the necessary operations of adding, deleting, and retaining words.
@@
We harvested VOA Special English vocabulary comprising 1,517 unique words (VOA1500).77endnote: 7We included VOA Special English Word Book Sections A-Z, Science Programs, and Organs of the Body hosted on Wikipedia (https://simple.wikipedia.org/wiki/Wikipedia:VOA_Special_English_Word_Book). We calculate the ratio of words that appear in the VOA1500 to those that do not, then report the natural logarithm of this ratio for each generated sample. Values above 0 indicate that the text contains more Special English words than non-Special English words, and a higher value indicates a greater presence of “easy” words.
@@ 5.2.4 Token Distribution Shift Analysis
To understand the impact of optimization guided by different rewards, we analyzed the shifts in token distribution in generations from reinforcement learning models compared to those generated by the supervised fine-tuning model from which they were initiated. Following the approach of Lin et al. (2023), we first generated a response using a model trained with reinforcement learning for a given query. We then used the supervised fine-tuning model to predict the most probable token at each position based on the context up to that point.
@@ 6.1 Quantitative Assessment
Table 2: Comparison of Gemma-2B’s performance across three approaches: the supervised fine-tuned baseline (SFT), reinforcement learning guided by ARI (using an intermediate checkpoint before significant policy gradient instability was observed, RLARI), and reinforcement learning guided by two accessibility measures (RLAM). SFT was fine-tuned using the Scientific Abstract-Significance Statement (SASS) corpus reported in Wang and Clark (2024). The columns labeled $\beta_{\text{WA}}$ and $\beta_{\text{SL}}$ pertain specifically to RLAM, where the rewards for average word accessibility and sentence length are balanced. The inference on the test split from SASS uses multinomial sampling. Metrics ARI, F-K, SARI, VOA, SL, WA, WL, and BS stand for Automated Readability Index, Flesch-Kincaid readability test, log ratio of VOA1500 vocabulary, sentence length, word accessibility, word length, and BERTScore (F1), respectively. Measures followed by a down arrow symbol ($\downarrow$) indicate that lower values are better. Numeric values in parentheses are the corresponding standard deviations. A paired two-tailed t-test was performed on observations of each measure between each model and the original abstracts. At a model-wise p-value of 0.05, measures that differ significantly from the SFT baseline are marked with an asterisk.
@@
We observe that reinforcement learning models trained with different rewards exhibit a notable reduction in reading level, bringing abstracts down to high school levels. The model directly guided by ARI (RLARI) achieves an ARI of 12.6, while the most performant model guided by accessibility measures (RLAM, $\beta_{\text{SL}}=4.0$ and $\beta_{\text{WA}}=0.2$) reaches 12.5, both aligning with the readability level expected for individuals who have completed K–12 education (approximately ARI 13). However, RLARI and RLAM models achieve these readability improvements in different ways. For RLAM models, better readability is achieved mostly through improved token-level accessibility. RLAM models show an increase in word accessibility from 0.5 to 0.8 compared to the supervised baseline. This increase in the natural logarithm suggests that words generated by RLAM models are, on average, 1.6 to 2.2 times more frequent in the English Wikipedia corpus than those generated by the SFT model. In comparison, RLARI’s 0.1-unit increase in word accessibility, although observed, does not result in a statistically significant change in word frequency compared to the SFT model. Similarly, the log ratio of the VOA1500 vocabulary in the RLAM models shows a significant improvement, with log ratios ranging from --0.02 to 0.08. This implies that for every 100 non-VOA1500 (or more complex) words generated, RLAM models can produce approximately 98 to 106 VOA1500 basic words. In contrast, the SFT and RLARI models exhibit VOA log ratios of --0.26 and --0.17, respectively, indicating that for every 100 non-VOA1500 words generated, these models produce only around 77 (84) VOA1500 words. The average word length in characters for RLAM models ranges from 4.8 to 4.9, slightly shorter than RLARI’s and outperforming SFT. Overall, the above evidence suggests that RLAM models achieve better readability by using more common, simpler, and shorter words.
@@ 6.2 Qualitative Analysis
We also found that the proportion of these artifacts steadily increases as training continues, and they are usually found in only a subset of samples across different experiments. Further, when $\beta_{\text{SL}}$ is set to a larger value, such as 0.5, the generations tend to cheat by generating readability-deflating trailers within the first few hundred iterations of optimization. To address this problem, we experimented with an additional heuristic—the standard deviation of sentence-level word accessibility. When this measure exceeds approximately 0.6, it indicates that readability-deflating trailers are beginning to accumulate. We had limited success using this extra reward and hypothesized that it might be too indirect for the language model to effectively learn from amid other signals that may covary with this measure. During manual examinations across experiments, no repetition was observed, which we had previously identified as a major concern in the SFT models (Wang and Clark, 2024).

### v2

@@ 4.1 Language Modeling via Proximal Policy Optimization
At the core of our approach is language modeling with Proximal Policy Optimization (PPO) (Schulman et al.,, 2017) guided by two accessibility measures. A causal language model trained on large corpora can generate the next token based on the current sequence, which is useful in the context of reinforcement learning for developing a policy model that determines the most appropriate next token to maximize the expected return in terms of document readability.
@@
Here, $J(\pi_{\theta})$ represents the expected return when following policy $\pi_{\theta}$. The reward $r(s_{t},a_{t})$ is estimated for each time step $t$ in the trajectory $\tau=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{T-1},a_{T-1})$, where $s_{t}$ is the sequence of tokens at time $t$, formed as the concatenation of $s_{0}$ and the tokens $a_{0},a_{1},\ldots,a_{t-1}$. This formula iteratively computes the rewards given the current sequence $s_{t}$ and the token $a_{t}$ chosen by the policy. The $\beta_{\text{KL}}$-weighted KL divergence term $\text{KL}(\pi_{\theta}(a_{t}\mid s_{t})\parallel\pi_{\theta_{\text{SFT}}}(a_{t}\mid s_{t}))$ is applied at every step of sequence generation to ensure the policy does not deviate significantly from the supervised fine-tuned model. This is crucial because, without such a constraint, the policy model might quickly learn to output whatever the reward model favors to maximize its return, which can lead to undesirable behaviors. For instance, the model might repeatedly output a frequent word (“is is is …”), achieving a high reward based solely on accessibility measures but lacking meaningful content. Following Stiennon et al., (2020), $\beta_{\text{KL}}$ is dynamically adjusted by targeting a specific KL divergence between $\pi_{\theta}$ and $\pi_{\theta_{\text{SFT}}}$ using a capped proportional controller in logarithmic space. The benefit of using a dynamic KL control, as opposed to a fixed one, is that it allows the model to adapt more flexibly to different stages of training, accommodating varying levels of KL divergence between the policy and SFT model. For further details, see Appendix A for the specific PPO implementation we are using and Appendix B for the dynamic proportional controller.
@@ 4.2 Reward Function
The reward function evaluates the overall quality of the output ($s_{T}$). After the initial failures of testing a traditional readability measure (i.e., ARI) as the criterion, we decided to use a balance of two accessibility measures: average sentence length in words and word accessibility, adopted from Riddell and Igarashi, (2021), to guide the optimization.
@@
A word’s accessibility is approximated by how frequently it appears in a large reference corpus. We chose the English Wikipedia corpus due to its domain similarity and applied a Moses tokenizer, yielding a vocabulary of 14.6 million types from a total of 3.6 billion tokens. If a token is among the most common 100,000 types, we report its frequency per billion tokens as its accessibility measure. Otherwise, we estimate its frequency using ridge regression with an $\ell 2$-norm coefficient equal to $1.0$. This model allows us to make serviceable estimates of the frequency of arbitrary tokens, including tokens that do not appear in the reference corpus. This model takes as input the token’s length in Unicode code points, its byte unigrams, byte bigrams, and byte trigrams. The model estimates the token’s log frequency per 1 billion tokens. We used the natural logarithm of frequencies per billion tokens as the measure of word accessibility. [^3] We have faithfully followed the experiment of Riddell and Igarashi, (2021) with three differences. First, our reference corpus is the English Wikipedia, whereas the original study used the Common Crawl News corpus. Second, we did not discard duplicated sentences as Riddell and Igarashi, (2021) did, because we found that sentence duplication is not common in Wikipedia. Third, the original study reported word inaccessibility scores by negating the logarithm of frequency per billion. We report accessibility, without negation, because it is more naturally suited to serve as a reward. Refer to Riddell and Igarashi, (2021, pp. 1186–1187) for the training of the ridge regression. For example, the accessibility score for “big” is 11.8, while “colossal” scores 7.3. Despite being comparable in meaning, the model’s production of the latter will receive fewer rewards. Coefficient $\beta_{\text{{wa}}}$ is to control the scale of the credit given for word accessibility.
@@
Sentence length is also determined by a Moses tokenizer, which preserves hyphenation and splits contractions. [^4] For the Moses rule-based tokenizer, we use the sacremoses Python package. We negate the value of sentence length for intuitive calculation of the rewards for optimization.
@@ 5.1 Training
We initialized the policy models $\pi_{\theta}$ by adopting the Gemma-2B checkpoint reported by Wang and Clark, (2025) ($\pi_{\theta_{\text{SFT}}}$). The original Gemma-2B was trained on three trillion tokens, consisting of publicly available data as well as proprietary datasets comprising “primarily English data from web documents” (Mesnard et al.,, 2024). The specific checkpoint we adopted was fine-tuned using the SASS corpus in a straightforward manner. It was chosen for its strong performance in simplification quality, its faithfulness, and its relatively compact size.
@@
The two accessibility rewards were weighted as follows: the word accessibility reward was set to $\beta_{\text{WA}}=4.0$, and we report four models with gradually increasing $\beta_{\text{SL}}$ values of $0.05$, $0.08$, $0.1$, and $0.2$. [^5] Selecting hyperparameters for PPO training is a computationally intensive task, particularly due to the sensitivity of reward model weights. We grounded our search in intuitive defaults commonly adopted in PPO literature—learning rate, batch size, and value function coefficient—and verified training stability on a surrogate task (IMDb sentiment alteration). After confirming reliable PPO dynamics, we conducted a targeted random sweep over $\beta_{WA}$ and found that values near $\beta_{WA}=4.0$ consistently produced stable learning and favorable trade-offs between readability and semantic retention. We then refined the joint configuration of $\beta_{KL}$ and $\beta_{WA}$ to further stabilize optimization and prevent collapse and semantic drift. While we acknowledge that this hyperparameter configuration may not be globally optimal, it reflects a deliberate balance between simplicity and effectiveness. Rather than relying on a separate language model trained with extensive human annotations to approximate human evaluation (Ouyang et al.,, 2022), our reward function—implemented in just a few lines of Python—offers an elegant and computationally efficient solution for guiding the simplification process. In addition, because word accessibility is in logarithmic space, we subtracted 10 from the estimated word accessibility and reset any values lower than 10 to 0 to keep them within a reasonable range. For adaptive control of the per-token semantic reward, we started with an initial $\beta_{\text{KL}}=0.2$ and targeted a KL divergence of $8.0$ nats during the training course, capping it in the range between $0.15$ and $0.25$. We used a micro batch size of $4$, with each sequence used to run the PPO algorithm for $4$ epochs using importance sampling, with gradient accumulation steps set to $4$. We used a clip range of $0.2$ for the policy gradient and value function estimation to ensure stability. The value function coefficient was set to $0.1$. The optimization used standard AdamW optimizer parameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1\times 10^{-8}$) (Kingma and Ba,, 2015; Loshchilov and Hutter,, 2019). The learning rate was fixed at $1\times 10^{-6}$. The training was conducted on 2 H100 (80GB) GPUs using mixed precision training in bfloat16. The sampling temperature was set to $0.7$ in the rollout phase. Following Huang et al., (2024), we assigned unfinished roll-outs (indicated by the missing end of sequence token) with a fixed low score to encourage the model to generate complete simplified narratives. We also report a model trained as dictated by ARI [^6] We began with pilot experiments using ARI as the sole reward, and observed that it encouraged the model to take shortcuts, often collapsing outputs into overly compressed summaries—sometimes as short as a single sentence. While such ultra-brief abstracts resemble the desired behavior in extreme summarization tasks (Narayan et al.,, 2018), they proved suboptimal for enhancing accessibility. These results underscored the need to explicitly guard against excessive sentence shortening; without such constraints, reinforcement learning risks compromising the semantic fidelity of the output. (i.e., RLARI; described in Section 1.2), with all other parameters kept the same as those guided by the accessibility measures.
@@ 5.2.1 Semantic Retention
BERTScore calculates the cosine similarity between each token in the candidate sentence and each token in the reference sentence using contextual embeddings from a pre-trained language model; its results align well with human judgment on semantic similarity evaluation (Zhang et al.,, 2020). It is not directly influenced by lexical overlap, making it more suitable for evaluating simplification systems than are metrics that rely on matching words, such as BLEU (Papineni et al.,, 2002). For our evaluation, we used embeddings from the 18th layer of a BERT-large-uncased model and reported the F1 score. This choice is based on prior findings indicating that the 18th layer yields a strong Pearson correlation (0.72) on the WMT16 To-English benchmark (Zhang et al.,, 2020).
@@ 5.2.2 Simplification & Accessibility
SARI (System output Against References and against the Input sentence) is specifically designed to evaluate text simplification (Xu et al.,, 2016). It aims to measure how well a simplified text retains the original meaning while improving readability. SARI provides a balanced measure of how well a text simplification system performs by focusing on the necessary operations of adding, deleting, and retaining words.
@@
We harvested VOA Special English vocabulary comprising 1,517 unique words (VOA1500). [^7] We included VOA Special English Word Book Sections A-Z, Science Programs, and Organs of the Body hosted on Wikipedia (https://simple.wikipedia.org/wiki/Wikipedia:VOA_Special_English_Word_Book). We calculate the ratio of words that appear in the VOA1500 to those that do not, then report the natural logarithm of this ratio for each generated sample. Values above 0 indicate that the text contains more Special English words than non-Special English words, and a higher value indicates a greater presence of “easy” words. We chose the VOA word list because it is a well-known, publicly available vocabulary list specifically designed for learners of English as a second language. Its use aligns with established heuristics in readability research, which suggest that texts using learner-friendly words tend to be more accessible.
@@ 5.2.4 Token Distribution Shift Analysis
To understand the impact of optimization guided by different rewards, we analyzed the shifts in token distribution in generations from reinforcement learning models compared to those generated by the supervised fine-tuning model from which they were initiated. Following the approach of Lin et al., (2023), we first generated a response using a model trained with reinforcement learning for a given query. We then used the supervised fine-tuning model to predict the most probable token at each position based on the context up to that point.
@@ 6.1 Quantitative Assessment
Table 2: Comparison of Gemma-2B’s performance across three approaches: the supervised fine-tuned baseline (SFT), reinforcement learning guided by ARI (using an intermediate checkpoint before significant policy gradient instability was observed, RLARI), and reinforcement learning guided by two accessibility measures (RLAM). SFT was fine-tuned using the Scientific Abstract-Significance Statement (SASS) corpus reported in Wang and Clark, (2025). The columns labeled $\beta_{\text{WA}}$ and $\beta_{\text{SL}}$ pertain specifically to RLAM, where the rewards for average word accessibility and sentence length are balanced. The inference on the test split from SASS uses multinomial sampling. Metrics ARI, F-K, SARI, VOA, SL, WA, WL, and BS stand for Automated Readability Index, Flesch-Kincaid readability test, log ratio of VOA1500 vocabulary, sentence length, word accessibility, word length, and BERTScore (F1), respectively. Measures followed by a down arrow symbol ($\downarrow$) indicate that lower values are better. Numeric values in parentheses are the corresponding standard deviations. A paired two-tailed t-test was performed on observations of each measure between each model and the original abstracts. At a model-wise p-value of 0.05, measures that differ significantly from the SFT baseline are marked with an asterisk.
@@
We observe that reinforcement learning models trained with different rewards exhibit a notable reduction in reading level, bringing abstracts down to high school levels. The model directly guided by ARI (RLARI) achieves an ARI of 12.6, while the most performant model guided by accessibility measures (RLAM, $\beta_{\text{SL}}=4.0$ and $\beta_{\text{WA}}=0.2$) reaches 12.5, both aligning with the readability level expected for individuals who have completed K-12 education (approximately ARI 13). However, RLARI and RLAM models achieve these readability improvements in different ways. For RLAM models, better readability is achieved mostly through improved token-level accessibility. RLAM models show an increase in word accessibility from 0.5 to 0.8 compared to the supervised baseline. This increase in the natural logarithm suggests that words generated by RLAM models are, on average, 1.6 to 2.2 times more frequent in the English Wikipedia corpus than those generated by the SFT model. In comparison, RLARI’s 0.1-unit increase in word accessibility, although observed, does not result in a statistically significant change in word frequency compared to the SFT model. Similarly, the log ratio of the VOA1500 vocabulary in the RLAM models shows a significant improvement, with log ratios ranging from --0.02 to 0.08. This implies that for every 100 non-VOA1500 (or more complex) words generated, RLAM models can produce approximately 98 to 106 VOA1500 basic words. In contrast, the SFT and RLARI models exhibit VOA log ratios of --0.26 and --0.17, respectively, indicating that for every 100 non-VOA1500 words generated, these models produce only around 77 (84) VOA1500 words. The average word length in characters for RLAM models ranges from 4.8 to 4.9, slightly shorter than RLARI’s and outperforming SFT. Overall, the above evidence suggests that RLAM models achieve better readability by using more common, simpler, and shorter words.
@@ 6.2 Qualitative Analysis
We also found that the proportion of these artifacts steadily increases as training continues, and they are usually found in only a subset of samples across different experiments. Further, when $\beta_{\text{SL}}$ is set to a larger value, such as 0.5, the generations tend to cheat by generating readability-deflating trailers within the first few hundred iterations of optimization. To address this problem, we experimented with an additional heuristic—the standard deviation of sentence-level word accessibility. When this measure exceeds approximately 0.6, it indicates that readability-deflating trailers are beginning to accumulate. We had limited success using this extra reward and hypothesized that it might be too indirect for the language model to effectively learn from amid other signals that may covary with this measure. During manual examinations across experiments, no repetition was observed, which we had previously identified as a major concern in the SFT models (Wang and Clark,, 2025).